OpenStudy (anonymous):

How do I integrate 8sin(x)(cos(cos x))dx ?

OpenStudy (lgbasallote):

is this read as \[\int 8\sin x \cos (\cos x)dx\]

OpenStudy (lgbasallote):

interesting... let u = cos x du = -sin x \[\implies -8 \int \cos u du\] that should be easier now

OpenStudy (anonymous):

Thank you. that does make sense. It just confused me because of the cosine function! So -8sin(u) + C. I did that in my head, so sorry if there's some mistakes. But yes, the cosine function threw me off.

OpenStudy (lgbasallote):

dont forget to change u back into cos x;)

OpenStudy (anonymous):

Thank you so much! And I wont. :)

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (turingtest):

that integral almost breaks the wolf lol http://www.wolframalpha.com/input/?t=crmtb01&_=1343799917386&i=integral+8sinx*cos(cosx)dx&fp=1&incTime=true ^^^this only worked with extra computation time

OpenStudy (anonymous):

Yes, it timed out when I tried it, lol